Linux Consultants HOWTO

Martin Michlmayr, tbm@sypher.com, http://www.sypher.com/tbm v2.40, 7 December 1996 This document contains a listing of companies providing commercial Linux related support. It is maintained by Martin Michlmayr <tbm@sypher.com>.

Contents

1 About this document 6

2 Copyright Information 6

3 Argentina 7 3.1 CIBER*NET . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

4 Australia 8 4.1 ACAY Network Computing Pty Ltd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 4.2 Interactive Infotainment Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 4.3 InterWeft . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 4.4 IPAX Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

5 Austria (Österreich) 10 5.1 Bergmann Engineering and Consulting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 5.2 EDV-Consulting Wolfgang DREYER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 5.3 g.a.m.s. edv-dienstleistungen gmbh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 5.4 EDV-Beratung Friedrich Ganter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 5.5 Horus GmbH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 5.6 Alexander List . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

6 Belgium 14 6.1 Double Barrel Consultancy and Productions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

7 Canada 15 7.1 Cabala Systems Corp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 7.2 Dataden Computer Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 7.3 Datasoft Communications . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 7.4 Dynamicro Consulting Limited . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 7.5 Global Proximity Corporation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 7.6 G.M. Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 7.7 Interactive Electronic Design Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 7.8 Jones Computer Supplies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 7.9 MostlyLinux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 7.10 Gary Lawrence Murphy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 7.11 Ken Roberts Computer Consultants Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 7.12 Silver Technologies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

8 Europe 23 8.1 Jochen Hayek . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

9 Finland 24 9.1 Netsol Network Solutions Oy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

10 Germany (Deutschland) 25 10.1 Gerd Aschemann . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 10.2 Hans-Joachim Baader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 10.3 BEE Baastrup EDV-Entwicklung GmbH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 10.4 Jens Doll . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 10.5 EDV-Beratung Gert Döring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 10.6 Karl Feichtinger information systems & communications consulting . . . . . . . . . . . . . . . 28 10.7 Mike Gärtner . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 10.8 Gärtner Datensysteme . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 10.9 Ing. Büro Gerichhausen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 10.10IKS GmbH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 10.11LEMIS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 10.12MetaWorks GmbH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 10.13MU Softwareentwicklung . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 10.14B. Neubert EDV-Dienstleistungen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 10.15noris network GmbH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 10.16Ntakpe Management & Consulting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 10.17Martin Schulze . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 10.18Julian Howard Stacey (BSc Hons, Computers and Cybernetics) . . . . . . . . . . . . . . . . . 36 10.19S.u.S.E. Gmbh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 10.20Jörg Weber . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 10.21Arne Wichmann . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 10.22Michael Will Linux Systeme . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 10.23Bernd Wolf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 10.24WorNet Internetdienste . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

11 Israel 42 11.1 TclTek Ltd. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42

12 Italy 43 12.1 PROFUSO di Zanetti Giuseppe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 12.2 SIR s.r.l. - PublinetWork . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 12.3 Softmedia Spa . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

13 Japan 45 13.1 SoftWare Research Co. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

14 Mexico 45 14.1 Miguel de Icaza . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 14.2 Laboratorio de Sistemas Linux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

15 The Netherlands 47 15.1 Bitbybit Information Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 15.2 BitWizard B.V. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 15.3 Cistron Internet Services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 15.4 MANTRA Research . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

16 Norway 50 16.1 it-analyse as . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 16.2 NorConnect Internet Services AS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

17 Pakistan 51 17.1 Vision Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 18 Poland 52

18.1 Polskie Sieci Teleinformatyczne S.A. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

19 South Africa 53 19.1 Obsidian Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53

20 Sweden 53 20.1 Guide Datakonsult . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 20.2 Signum Support AB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 20.3 Tarkus AB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

21 Turkey 55 21.1 Hakan Kalyoncu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

22 United Kingdom (UK) 56 22.1 Cygnet Internet Services Ltd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 22.2 Paul Dunne . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 22.3 Laboratory for Micro Enterprise . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 22.4 Nene Information Sytems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

23 United States of America (USA) 59 23.1 Aegis Information Systems, Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 23.2 Rahim Azizarab . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 23.3 BRT Technical Services Corporation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 23.4 Colorado Computer Consultants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 23.5 M. Cooper . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 23.6 Couvares Consulting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 23.7 Crynwr Software . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 23.8 Daylight Software . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 23.9 DESIGN SCIENCE LABS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 23.10DigiTech, Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 23.11Electronic Information Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 23.12Etc Services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 23.13Evolve Computer Solutions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 23.14Christopher J. Fearnley . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 23.15Granite Computing Solutions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 23.16Steve Hedlund . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 23.17I-Link, Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 23.18InfoMagic, Inc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 23.19Internet Robotics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 23.20Internet Systems and Services . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 23.21Isle-Net Telecommunications . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 23.22Matthew Kaylor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 23.23Kensho, LLC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 23.24Kracked Rock Komputing, LLC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 23.25Kymsoft Consulting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 23.26LANscape . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 23.27LOD Communications . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 23.28NDA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 23.29NET-Community . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 23.30Neu-Direct Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 23.31The Network Wizards . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 23.32Alvin Oga . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 23.33onShore, Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 23.34Practical Network Design . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 23.35Probe Technology Inc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 23.36PSA AnswerDesk – Pencom Systems Administration . . . . . . . . . . . . . . . . . . . . . . . 83 23.37RCT Design . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 23.38Reboot Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 23.39RG Consulting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 23.40Bill Rousseau . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 23.41Softcraft Impresa . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 23.42Sophic Consulting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 23.43Earl A. Stutes, Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 23.44Triangle Software Corporation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 23.45tummy.com, ltd. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 23.46Vennerable Consultants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 23.47Vertex Development . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 23.48Gregg Weber . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 23.49Jim Willette . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 23.50Worldwide Solutions, Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 23.51xprt Computer Consuting, Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 23.52Yggdrasil Computing, Inc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 23.53ZEI Software . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95

1 About this document

This is the Linux Consultants HOWTO. It is a listing of companies providing commercial Linux related support. It is maintained by Martin Michlmayr <tbm@sypher.com>. If you need to know more about the Linux Documentation Project or about Linux HOWTO’s, feel free to contact the supervisor Greg Hankins <gregh@sunsite.unc.edu>. Greg Hankins will post the listing to several national and international newsgroups on a monthly basis. In addition, the Consultants HOWTO is available via FTP at ftp://ftp.sypher.com/tbm/HOWTO or can be found on the World Wide Web at http://www.sypher.com/tbm/Consultants-HOWTO. New versions of the Consultants HOWTO are always placed at these two sites first, so please be sure to check if the copy you are reading is out of date! Companies providing Linux support are invited to fill out the following form and contact me at tbm@sypher.com.

     Name:
     Address:
     Phone:
     Fax:
     EMail:
     Url:
     Contact:
     Type of support (via phone, EMail, inhouse, training etc.):
     Sample prices:

2 Copyright Information

This HOWTO is Copyright \begin{matrix}c \\ ◯ \end{matrix} 1996 by Martin Michlmayr. A verbatim copy may be reproduced or distributed in any medium physical or electronic without permission of the author. Translations are similarly permitted without express permission if it includes a notice on who translated it.

Short quotes may be used without prior consent by the author. Derivative work and partial distributions of the Consultants HOWTO must be accompanied with either a verbatim copy of this file or a pointer to the verbatim copy. Commercial redistribution is allowed and encouraged; however, the author would like to be notified of any such distributions. In short, we wish to promote dissemination of this information through as many channels as possible. However, we do wish to retain copyright on the HOWTO documents, and would like to be notified of any plans to redistribute the HOWTOs. We further want that all information provided in the HOWTOs is disseminated. If you have questions, please contact Greg Hankins, the Linux HOWTO coordinator, at gregh@sunsite.unc.edu.

3 Argentina

3.1 CIBER*NET

          Derqui 47
          5501 Godoy Cruz
          Mendoza
          Argentina

Phone:

(54-61)-243883 Fax:

None EMail:

afernand@planet.losandes.com.ar Url:

None Contact:

Alejandro Fernandez Herrero Type of support:

General Unix/Linux consulting, networking, connectivity with other operating systems, support, porting and web pages development. Onsite training.

Sample prices:

Upon request. Last modified:

October 25, 1996.

4 Australia

4.1 ACAY Network Computing Pty Ltd

          Level 1, 10 Help Street
          Chatswood, NSW 2067
          Australia

Phone:

+61 2 9411 7340 Fax:

+61 2 9411 7340 EMail:

support@acay.com.au Url:

http://www.acay.com.au Contact:

Kris Kulasekera <kris@acay.com.au> Type of support:

N/A Sample prices:

AUS$60/hour. Last modified:

October 18, 1996.

4.2 Interactive Infotainment Systems

          Sydney, NSW
          Australia

Phone:

015-955-042 Fax:

None EMail:

anthony@rumble.waratah.id.au Url:

http://www.infotainment.com.au Contact:

Anthony Rumble Type of support:

Phone, EMail, inhouse, training, porting, development, web/CGI, networking. Sample prices:

Upon request. Last modified:

September 24, 1996.

4.3 InterWeft

     Melbourne
     Australia

Phone:

+61 (0)3 9735 3586 Fax:

+61 (0)3 9735 3586 EMail:

info@interweft.com.au Url:

http://www.interweft.com.au Contact:

Robert Hart Type of support:

Strategic IT business planning. Data/voice networking (design, specification and integration), Internet connectivity planning, design, installation, configuration and security. Linux/Unix installation, configuration, training and support. Author of Linux PPP-HOWTO and mail2news Mini-HOWTO

Sample prices:

Upon request. Quick help via EMail is provided free as long as the work load permits - particularly for Linux PPP.

Last modified:

November 30, 1996.

4.4 IPAX Systems

     GPO Box 94A
     Melbourne, Vic
     Australia 3001

Phone:

+61 3 9887 1984 Fax:

+61 3 9801 8533 EMail:

info@ipax.com.au Url:

http://www.ipax.com.au Contact:

Adam Neat Type of support:

We do System/Network design. We use Linux for the majority of the Unix machines we install. 24 hour NOC services are also available. Linux is used for this. We will provide free advice via EMail and fax.

Sample prices:

Free for advice. Last modified:

September 27, 1996.

5 Austria (Österreich)

5.1 Bergmann Engineering and Consulting

     Dirmhirngasse 27/118
     A-1230 Wien
     Austria

Phone:

+43 (1) 8695186, +43 (2254) 7802703 or +43 (664) 3573565 Fax:

+43 (1) 319956110 EMail:

office@bec.co.at Url:

http://www.bec.co.at Contact:

Josef Bergmann Type of support:

Phone, EMail and inhouse. Installation, troubleshooting, training, administration, consulting, Internet and Intranet connectivity, Linux file, print, fax, mail, web, FTP, DNS, cache, modem servers. Custom software development (C and C++) for scientific, industrial or business applications

Sample prices:

Base price ATS 950/hour (about US$80). Last modified:

October 9, 1996.

5.2 EDV-Consulting Wolfgang DREYER

     Hintnausdorf 14/1
     A-9300 St. Veit/Glan
     Austria

Phone:

+43 4212 60040 Fax:

+43 4212 60044 EMail:

support@quant-x.com Url:

http://www.quant-x.at Contact:

Fritz Ganter Type of support:

EMail, telephone, onsite. Sample prices:

30 days getting started US$150.

Last modified:

October 10, 1996.

5.3 g.a.m.s. edv-dienstleistungen gmbh

     Sch\"{o}nbrunnerstr. 179/II/4
     A-1120 Vienna
     Austria

Phone:

+431 8176230 Fax:

+431 8176230-17 EMail:

info@gams.at Url:

http://www.gams.at Contact:

Vinzenz Grabner, Michael Lausch, Rainer Hochreiter Type of support:

Phone or EMail for commercial customers. Linux installation and configuration, networking, security policies, firewall configurations. Programming of Internet applications (e.g. http://www.gams.at/ubahn), Internet service provider (leased lines, dial-up, WWW, ...) We are using the Red Hat Linux distribution.

Sample prices:

Upon request. Last modified:

October 9, 1996.

5.4 EDV-Beratung Friedrich Ganter

     Grazerstra{\ss}e 26a
     A-8045 Graz
     Austria

Phone:

+43 316 688112 or +43 664 1021621 Fax:

None EMail:

ganter@fvkma.tu-graz.ac.at Url:

http://fvkma.tu-graz.ac.at/ganter/ganter.html Contact:

Friedrich Ganter Type of support:

Phone, EMail, inhouse, training. Sample prices:

US$30-$60/hour. Last modified:

September 12, 1996.

5.5 Horus GmbH

          Jakob-Haringer-Stra{\ss}e 8
          A-5020 Salzburg
          Austria

Phone:

+43 (0)662 450363 Fax:

+43 (0)662 450364 EMail:

info@horus.com Url:

http://www.horus.com Contact:

Thomas ’Dune’ Freina Type of support:

EMail and inhouse. Consulting, installation, training and programming.

Sample prices:

Base price is ATS 1250 per hour (about US$115). Last modified:

October 16, 1996.

5.6 Alexander List

          Neue-Welt-H\"{o}he 52A
          A-8042 Graz
          Austria

Phone:

+43-316-474737 Fax:

None EMail:

alexlist@sbox.tu-graz.ac.at Url:

http://www.sbox.tu-graz.ac.at/home/alexlist Contact:

Alexander List Type of support:

Phone, EMail, inhouse, training. Specialized on ISDN solutions. Will also set up your LAN and clients running other operating systems if desired.

Sample prices:

Basic rate US$60/hour. To be negotiated, depending on type and duration of job. Last modified:

October 7, 1996.

6 Belgium

6.1 Double Barrel Consultancy and Productions

          Sportstraat 28
          9000 Gent
          Belgium

Phone:

32-9-2227764 Fax:

32-9-2224976 EMail:

mvergall@double-barrel.be Url:

http://www.double-barrel.be Contact:

Michael C. Vergallen Type of support:

Phone, EMail, inhouse. Sample prices:

1500 BEF/hour V.A.T. excl. (1 hour minimum). Last modified:

November 29, 1996.

7 Canada

7.1 Cabala Systems Corp.

          801 3760 Albert St.
          Burnaby, B.C.
          V5C 5Y8
          Canada

Phone:

(604) 294-1721 Fax:

(604) 294-1751 EMail:

kcarson@direct.ca Url:

None Contact:

Kevin Carson Type of support:

Phone, EMail, onsite, administration, training and programming. We cover all Linux related areas but specialities comprise of X11 graphics, security, networking, and 24/7/365 applications.

Sample prices:

Dependent on activity/contract. Last modified:

September 23, 1996.

7.2 Dataden Computer Systems

     895 2nd Avenue
     Ile Bizard, Quebec
     Canada, H9C 1K3

Phone:

(514) 891-2293 Fax:

(514) 696-0848 EMail:

arseneau@dataden.com Url:

http://www.dataden.com Contact:

Danny Arseneau Type of support:

Linux installation, configuration and administration. Installation, configuration and administration of Samba on Linux. Setting up Internet connections with Linux as a gateway. Setting up firewalls and other security systems with Linux. Setting up Linux as a network file and print server. Integration of Linux into existing networks. Software development for Unix in C/C++.

Sample prices:

CDN$60/hour. Negotiable for longer commitments. Last modified:

October 1, 1996.

7.3 Datasoft Communications

          36 Ellendale Dr.
          Scarborough, ON M1P 1N2
          Canada

Phone:

416-285-0803 Fax:

416-285-5607 EMail:

mark@datasoft.on.ca Url:

http://www.datasoft.on.ca Contact:

Mark Buckaway Type of support:

Linux support via phone, EMail and inhouse. Please see http://www.datasoft.on.ca/about.html Sample prices:

CAN$85/hour (1 hour minimum) or a flat rate contract (Please see

http://www.datasoft.on.ca/prices.html) Last modified:

November 30, 1996.

7.4 Dynamicro Consulting Limited

          282 Waratah Avenue
          Newmarket, Ont. L3Y 5J2
          Canada

Phone:

Available to clients. Fax:

Available to clients. EMail:

glouis@dynamicro.on.ca Url:

None Contact:

Greg Louis Type of support:

Phone, EMail, inhouse, training. Also custom software development for scientific, engineering, monitoring/control or business applications.

Sample prices:

CAN$125/hour for short-term contracts, negotiable for longer commitments. Last modified:

December 1, 1996.

7.5 Global Proximity Corporation

     959 Alice Street
     Woodstock
     Ontario N4S 8B1
     Canada

Phone:

519-421-3541 Fax:

519-421-2107 EMail:

CTyler@Oxford.net Url:

http://www.he.net/˜ globprox Contact:

Chris Tyler Type of support:

Full range of local and remote consulting services, ranging from installation support or the creation of small CGI scripts through the provision of full-service managed-Intranet services (including HW/SW/comm/distaster recovery planning/admin).

Sample prices:

Base price CAN$100/hour (about US$80), lower for large projects, higher for urgent work. Last modified:

September 12, 1996.

7.6 G.M. Systems

          2570 Seaview Road
          Mill Bay, BC. V0R 2P0
          Canada

Phone:

(250) 743-1500 Fax:

(250) 743-1500 EMail:

george@gmsys.com Url:

http://www.ultranet.ca/gmsys (Will become http://www.gmsys.com soon) Contact:

George A. Farris Type of support:

Phone, EMail, inhouse, training, equipment sales and setup with Linux pre-installed. Sample prices:

Contract dependent. Last modified:

October 5, 1996.

7.7 Interactive Electronic Design Inc.

          338 Westcourt Pl.
          Waterloo, ON, N2L 2R7
          Canada

Phone:

519-725-7353 Fax:

519-725-7353 EMail:

orders@ied.com Url:

None Contact:

Jan Vicherek Type of support:

Phone, EMail, inhouse, training, tutoring, Internet connections setup, custom programming. Sample prices:

CAN$20-$40/hour plus expenses. Last modified:

September 24, 1996.

7.8 Jones Computer Supplies

          RR \#3
          Perth, Ontario K7H 3C5
          Canada

Phone:

(613)267-6704 Fax:

(613)264-0372 EMail:

sos@jcs.on.ca Url:

http://www.jcs.on.ca Contact:

Peter Jones Type of support:

Installation of Linux, network configuration (firewalls, gateways, FTP, POP and WWW servers). Sample prices:

CAN$60-$160/hour depending on the work being performed. Last modified:

November 17, 1996.

7.9 MostlyLinux

     Calgary
     Alberta
     Canada

Phone:

(403) 209-9406 Fax:

(403) 285-1399 EMail:

info@mostlylinux.ab.ca Url:

http://www.mostlylinux.ab.ca Contact:

Scott Barker Type of support:

We do everything, from vanilla installs, to specific customizations and hacks, to ongoing system administration. However, we do not do telephone-only support. We will do telephone support only if we can have direct access to the system, such as through the Internet or through a no-charge dial-in. Also, since we are a very small company (at the moment), we will not return long-distance phone-calls, unless they are toll free.

Sample prices:

$20/hour (personal); $30/hour (commercial). Last modified:

November 9, 1996.

7.10 Gary Lawrence Murphy

     RR\#1 Box F3
     Sauble Beach, Ontario
     Canada

Phone:

519-422-1150 Fax:

519-422-2723 EMail:

garym@sos.on.ca Url:

http://www.sos.on.ca/˜ garym Contact:

Gary Lawrence Murphy Type of support:

Phone, EMail, telnet, direct or distance training (via MUD). Web site installations and design, Linux installations particularly for web sites and gateways. Irix and Solaris also supported. Projects of note: Ontario Science Centre Information Highway exhibit and OSC web site, Bell Canada ’AutoLinx’ intranets, SOS Internet Corp web site and ISP service, Town of Wiarton web site.

Sample prices:

CDN$60/hour but have been known to work for free or for barter for a right cause. Last modified:

October 17, 1996.

7.11 Ken Roberts Computer Consultants Inc.

     148 York Street, 2nd Floor
     London, Ontario
     N6A 1A9
     Canada

Phone:

(519) 672-8844 Fax:

(519) 672-3528 EMail:

krcc@mirror.org Url:

http://www.mirror.org/krcc.html Contact:

Ken Roberts Type of support:

Custom applications software, C programming, database applications including DBase and Appgen file formats. We have worked with Unix and Linux for seven years and have must-be-reliable applications operating on systems in over 50 Ontario communities. Familiar with networks. Remote programming, debugging, installation and support.

Sample prices:

CAN$120/hour. Last modified:

November 27, 1996.

7.12 Silver Technologies

     Box 22123
     Regina
     Saskatchewan S4S 7H4
     Canada

Phone:

306-596-1332 Fax:

306-585-2878 EMail:

rcp@dlcwest.com Url:

None Contact:

Richard Perrin Type of support:

System and network administration, onsite consultation, programming, Linux publications. Sample prices:

CDN$20-$50/hour. Last modified:

September 12, 1996.

8 Europe

Companies which are not bound to a particular country within Europe.

8.1 Jochen Hayek

          Augsburger Str. 33
          D-10789 Berlin
          Germany

Phone:

+49(171)350.28.73 (mobile), +49(30)214.798.80 Fax:

+49(30)214.798.81 EMail:

Jochen Hayek@ACM.org Url:

http://www.ACM.org/˜ Jochen Hayek Contact:

Jochen Hayek Type of support:

To be negotiated - quite unrestricted. Sample prices:

ECU 60/hour. Last modified:

November 5, 1996.

9 Finland

9.1 Netsol Network Solutions Oy

          Hiihtotie 3A
          FIN-01280 Vantaa
          Finland

Phone:

+358 50 5666898 Fax:

+358 9 34090404 EMail:

netsol@netsol.fi Url:

http://www.netsol.fi Contact:

Timo Virtaneva Type of support:

Inhouse, training. Support for Internet connections, firewalls and networking. Sample prices:

Upon request. Last modified:

September 15, 1996.

10 Germany (Deutschland)

10.1 Gerd Aschemann

          Osannstr. 49
          D-64285 Darmstadt
          Germany

Phone:

+49-6151-16-2259 Fax:

+49-6151-16-3052 EMail:

aschemann@Informatik.TH-Darmstadt.de Url:

http://www.informatik.th-darmstadt.de/˜ ascheman Contact:

Gerd Aschemann Type of support:

Phone, EMail, inhouse, training. Sample prices:

US$100/hour. Last modified:

October 8, 1996.

10.2 Hans-Joachim Baader

          Poppelfeld 9
          D-76646 Bruchsal
          Germany

Phone:

+49 7257 930142 (Mostly outside normal business hours) Fax:

+49 7257 930143 EMail:

hans@grumbeer.inka.de Url:

None Contact:

Hans-Joachim Baader Type of support:

Free support by EMail, phone, fax if I can answer by EMail. Linux/Unix programming, system administration, and consulting.

Sample prices:

Upon request. Last modified:

October 19, 1996.

10.3 BEE Baastrup EDV-Entwicklung GmbH

          Mallinkrodtstra{\ss}e 7
          D-44145 Dortmund
          Germany

Phone:

+49-231-9812060 Fax:

+49-231-9812062 EMail:

info@bee.de Url:

http://www.bee.de Contact:

Jörn Baastrup, Jürgen Stanke, Jörg Bellwinkel Type of support:

We’re concentrating in Linux and are writing software for customers. Different supports for Linux users: Internet, dial-in, ISDN-WAN...

Sample prices:

Take a look at our web site. Last modified:

October 7, 1996.

10.4 Jens Doll

          Hamburg
          Germany

Phone:

+49 40 67391235 Fax:

+49 40 67391237 EMail:

100117.3615@compuserve.com Url:

None Contact:

Jens Doll Type of support:

Inhouse, training. Sample prices:

Upon request. Last modified:

November 5, 1996.

10.5 EDV-Beratung Gert Döring

     Frankfurter Ring 24
     D-80807 M\"{u}nchen
     Germany

Phone:

+49 89 3545979 Fax:

+49 89 3545980 EMail:

gert@greenie.net Url:

None Contact:

Gert Döring, Simone Demmel Type of support:

Various (networking, LAN/WAN planning and implementation, system integration, system programming, training, general support).

Sample prices:

Prices vary depending on the job. Short jobs are at DM 95/hour, longer jobs at DM 700/day. Prices can be negotiated. Significantly lower prices for non-profit organizations.

Last modified:

October 19, 1996.

10.6 Karl Feichtinger information systems & communications consulting

     Ziegelh\"{u}tte 4
     D-66862 Kindsbach
     Germany

Phone:

+49 6371 60056 Fax:

+49 6371 69957 EMail:

feichtinger@iccc.se Url:

None Contact:

Dipl.-Inform. Karl Feichtinger Type of support:

Entwicklung und Administration von Internet und Intranet-Lösungen, Aufbau von EMail-Netzwerken, Installation und Wartung von Linux-Server-Systemen. Schulung, Support, Fernwartung.

Sample prices:

Tagessatz DM 1200, Festpreise nach Vereinbarung. Last modified:

October 22, 1996.

10.7 Mike Gärtner

     Am Harthwald 14
     D-09123 Chemnitz
     Germany

Phone:

+49-172-3526586 Fax:

+49-371-2600099, mailbox with support for Linux EMail:

mg@genyosha.in-chemnitz.de, mike@forum.org.kh (until January 1997) Url:

None Contact:

Mike Gärtner Type of support:

Phone (only in German, no price), EMail (only in German), inhouse (only in Chemnitz). Sample prices:

Upon request. Last modified:

October 13, 1996.

10.8 Gärtner Datensysteme

          Hamburger Str. 273a
          38114 Braunschweig
          Germany

Phone:

+49 531 2335555 Fax:

+49 531 2335556 EMail:

info@gaertner.de Url:

http://www.gaertner.de Type of support:

Phone, EMail, inhouse, training. Sample prices:

Depending on the job, starting at DM 1800 for a full day. Last modified:

September 13, 1996.

10.9 Ing. Büro Gerichhausen

          Steinsstra{\ss}e 103
          D-41199 M\"{o}nchengladbach
          Germany

Phone:

+49 2166 15068 Fax:

+49 2166 10159 EMail:

support@mail.niederrhein.de Url:

http://www.niederrhein.de Contact:

Christoph Adomeit Type of support:

Inhouse, training, phone, EMail, Internet. Sample prices:

DM 155/hour. Last modified:

September 12, 1996.

10.10 IKS GmbH

          Wildenbruchstra{\ss}e 15
          D-07745 Jena
          Germany

Phone:

+49 3641 675640 Fax:

+49 3641 675655 EMail:

info@iks-jena.de Url:

http://www.iks-jena.de Contact:

Erik Heinz Type of support:

We offer phone/EMail support, installation, remote administration, training. We also do network planning and setup, programming, and provide full Internet service.

Sample prices:

DM 150/hour for consulting. Contract prices available. Last modified:

November 5, 1996.

10.11 LEMIS

          Schellnhausen 2
          D-36325 Feldatal
          Germany

Phone:

+49-6637-919123 Fax:

+49-6637-919122 EMail:

lemis@lemis.de Url:

None Contact:

Greg Lehey Type of support:

Phone, EMail, fax, inhouse, training. Sample prices:

Typically US$100 or DM 150/hour. Significant reductions for non-profit organizations. Last modified:

September 23, 1996.

10.12 MetaWorks GmbH

          Florinsmarkt 5
          D-56068 Koblenz
          Germany

Phone:

+49 261 37773 Fax:

+49 261 37775 EMail:

support@metaworks.de Url:

http://www.metaworks.de Contact:

Andreas Heilwagen Type of support:

Phone, (EMail), inhouse, training. Sample prices:

Upon request. Last modified:

October 14, 1996.

10.13 MU Softwareentwicklung

          Wacholderweg 14
          D-70597 Stuttgart
          Germany

Phone:

+49-711-7676184 Fax:

+49-711-7676185 EMail:

info@musoftware.com Url:

http://www.musoftware.com Contact:

Ullrich von Bassewitz <uz@musoftware.com> Type of support:

Phone, EMail, inhouse, training (as needed). Sample prices:

Upon request. Last modified:

October 22, 1996.

10.14 B. Neubert EDV-Dienstleistungen

     Parkweg 16
     Scheidt
     D-66133 Saarbr\"{u}cken
     Germany

Phone:

0681-817629 Fax:

Upon request. EMail:

neubert@lusi.uni-sb.de (bevorzugt/preferred) Url:

http://www.uni-sb.de/˜ ph11rsbn Contact:

Boris Neubert Type of support (German):

Die Beratungstätigkeit umfaßt Hilfestellung bei Hardware- und Software-Kauf sowie Installation von Hardwarekomponenten und Anwenderprogrammen, Netzwerklösungen im Unix-Bereich, Vermittlung von Netzzugängen, sowie die allgemeine Informationsbeschaffung im wissenschaftlichen Bereich und verwandte Dienstleistungen.

Type of support (English):

Consulting comprises assistance with hardware and software purchase, installation of hardware components and user programs, networking solutions under Unix, mediation of network access, general information retrieval in the scientific field and related services.

Sample prices:

DM 40/hour (zuzüglich Spesen/plus expenses). Last modified:

September 13, 1996.

10.15 noris network GmbH

     Schleiermacherstrasss 12
     D-90491 N\"{u}rnberg
     Germany

Phone:

+49 911 59818 0 Fax:

+49 911 59818 11 EMail:

info@noris.de Url:

http://www.noris.de Contact:

Matthias Urlichs <smurf@noris.de> Type of support:

Remote login, phone, EMail, inhouse. Sample prices:

DM 150/hour for phone or inhouse support. Last modified:

September 17, 1996.

10.16 Ntakpe Management & Consulting

          Waldweg 9
          D-85386 Eching/Munich
          Germany

Phone:

+49 89 327155-0 Fax:

+49 89 327155-99 EMail:

ntakpe@ntakpe.com Url:

http://www.ntakpe.com Contact:

Jean Louis Ntakpe Type of support:

Phone, EMail, inhouse training, project consulting.

Sample prices:

Upon request. Last modified:

September 29, 1996.

10.17 Martin Schulze

          Artillerieweg 55a
          D-26129 Oldenburg
          Germany

Phone:

+49-441-7780584 Fax:

+49-441-777884 EMail:

joey@troi.orgatech.de Url:

http://home.pages.de/˜ joey Contact:

Martin Schulze Type of support:

Phone, EMail, inhouse, training. As needed. Sample prices:

Upon request. Last modified:

October 12, 1996.

10.18 Julian Howard Stacey (BSc Hons, Computers and Cybernetics)

          Holz Stra{\ss}e 27d
          D-80469 Munich
          Germany

Phone:

+49 89 268616 Fax:

+49 89 2608126 EMail:

jhs@freebsd.org Url:

http://www.freebsd.org/˜ jhs/vsl Contact:

Julian Stacey Type of support:

Phone (only after prepaid contract), EMail, training at customer site custom development on and offsite, system config, system administration, etc. Mainly FreeBSD based, but also Linux etc. I provide consulting in Bayern (Germany) and South East England.

Sample prices:

Upon request. Last modified:

October 11, 1996.

10.19 S.u.S.E. Gmbh

     Gebhardtstr. 2.
     D-90762 F\"{u}rth
     Germany

Phone:

+49 911 32471-56 or +49 911 7405331 Fax:

+49 911 3206727 EMail:

psupport@suse.de Url:

http://www.suse.de Contact:

Klaus Franken, Bodo Bauer Type of support:

Support mit/ohne Wartungsverträge, Entwicklung, WWW-Applikationen, Internet- und Intranet-Anbindung (z.B: ISDN), spezielle Hardwareanbindung, Beratung.

Sample prices:

          Supportvertrag
                  DM 240/Monat
                          Beinhaltet Anfragen per EMail und per Fax f\"{u}r zwei
                          Ansprechpartner, die sich ohne langwierige Forschungs- und
                          Entwicklungsarbeit beantworten l\"{a}{\ss}t.
          Ohne Vertrag
                  DM 180/Stunde = DM 15/5 Minuten
          Supportvertrag, Zusatzdienstleistung
                  DM 144/Stunde = DM 12/5 Minuten

Last modified:

October 14, 1996.

10.20 Jörg Weber

          Sursstr.1
          D-45143 Essen
          Germany

Phone:

+49-172-2186328 Fax:

+49-201-312677 EMail:

kufi@kufis.ruhr.de Url:

None Contact:

Jörg Weber Type of support:

Training, remote login, phone, EMail or by myself. System installation and maintainance (Standalone, LAN, WAN). Trouble shooting.

Sample prices:

Free for privates and non-profit organizations if it doesn’t take too much time. In general DM 120/hour plus expenses and VAT.

Last modified:

October 14, 1996.

10.21 Arne Wichmann

     Kaiserstra{\ss}e 35
     D-66121 Saarbruecken
     Germany

Phone:

Upon request. Fax:

None EMail:

aw@math.uni-sb.de Url:

http://vieta.math.uni-sb.de/˜ aw Contact:

Arne Wichmann Type of support:

Training, network purchase assistance, installation and support. Sample prices:

DM 40/hour. Last modified:

October 17, 1996.

10. Germany (Deutschland) 40

10.22 Michael Will Linux Systeme

Fichtenweg 1/1
D-72076 T\"{u}bingen
Germany

Phone:

07071-969063 (+49-7071-969063 for international calls)

Fax:

07071-969064

EMail:

Michael.Will@student.uni-tuebingen.de

Url:

http://www-ti.uni-tuebingen.de/˜ will

Contact:

Michael Will

Type of support:

Various

Sample prices:

These are only samples and will be adjusted where acceptable. In general, I do free support for home users and universities with light weight trouble shooting and consulting. When I have to invest more time, I charge money.

Consulting (hardware):
Home-users: free of charge
Universities: free of charge
Commercial: DM 80/hour
Consulting (coporate networking):
Home-users: free of charge
Universities: free of charge
Commercial: DM 120/hour
Trouble shooting (complex solutions):
Home-users: free of charge, but will only solve
        simple things.
Commercial: DM 80/hour
Universities: special contract (Werksvertr\"{a}ge)
Installation (hardware/software):
Home-users: DM 25/hour
Commercial: DM 120/hour
Universities: special contract (Werksvertr\"{a}ge)

Last modified:

October 1, 1996.

10.23 Bernd Wolf

     M\"{o}llner Str. 11, Wh. 3, Zi. 45a,
     D-18109 Rostock
     Germany

Phone:

0177/2350472 Fax:

None EMail:

wolf@egd.igd.fhg.de Url:

http://www.egd.igd.fhg.de/˜ wolf Contact:

Bernd Wolf Type of support:

EMail, eventuell direkt (talk), Phone. Verschiedenes, z.B.: System Installation/Administration, ”trouble shooting”, Hard-, Software- und Netz-Tips, Hilfe: learning by doing, ausserhalb Rostock: kommerziell (bis auf EMail) - nach Absprache, größere Aufträge.

Sample prices:

Preis nach Vereinbarung (Kommerziell), frei (andere). Last modified:

October 23, 1996.

10.24 WorNet Internetdienste

     Weidacher Hauptstr. 61
     D-82515 Wolfratshausen
     Germany

Phone:

+49 8171 22937 Fax:

+49 8171 22938 EMail:

wornet@wor.de Url:

http://www.wor.de Contact:

Christian Eich Type of support:

Phone, EMail, inhouse. Sample prices:

DM 80/hour. Last modified:

October 16, 1996.

11 Israel

11.1 TclTek Ltd.

          Mehalke HaMayim 12
          Jerusalem 91112
          Israel

Phone:

972.52.670.353 Fax:

972.2.563.0948 EMail:

benavrhm@tcltek.co.il Url:

http://www.tcltek.co.il Contact:

Jonathan Ben-Avraham Type of support:

Custom firewalls and routers, secure servers. Sample prices:

US$50/hour for companies, free via linux-il@linux.org.il Last modified:

October 31, 1996.

12 Italy

12.1 PROFUSO di Zanetti Giuseppe

     Please ask for the new address
     I-35031 Abano Terme (Padova)
     Italy

Phone:

++39 348 2220811 Fax:

None EMail:

profuso@profuso.com Url:

http://www.profuso.com/PROFUSO Contact:

Giuseppe Zanetti <beppe@iperv.it> Type of support:

Located in Veneto. Support Intel and Alpha Linux systems. System Integration and LAN technologies. Intranet. We offer Linux installation and Linux based solutions. EMail or phone support is available.

Sample prices:

Upon request. Last modified:

December 2, 1996.

12.2 SIR s.r.l. - PublinetWork

     P.zza Tavarone 2/22
     I-16123 Genova
     Italy

Phone:

+39 10 2463638 Fax:

+39 10 206761 EMail:

info@publinetwork.com Url:

http://www.publinetwork.com Contact:

Luca Maranzano Type of support:

Phone, EMail, remote login, telnet, inhouse (when possible), training, Internet and Intranet connectivity, web publishing, ready to run Linux servers with installation and configuration.

Sample prices:

Upon request. Last modified:

October 11, 1996.

12.3 Softmedia Spa

          via Pietrasanta 14
          I-Milano
          Italy

Phone:

(+39) 335-417677 Fax:

(+39) 2- 5730.7652 EMail:

venta@softmedia.it Url:

http://www.softmedia.it Contact:

Paolo Ventafridda Type of support:

Phone, inhouse, training and C programming. Sample prices:

Upon request. Last modified:

September 17, 1996.

13 Japan

13.1 SoftWare Research Co.

          Hayabusacho 2-11
          Chiyodaku, Tokyo
          Japan 102

Phone:

+81 3 3264 6171 Fax:

+81 3 3264 0647 EMail:

info@src.co.jp Url:

http://www.iijnet.or.jp/src, http://www.src.co.jp Contact:

Hayama Yossihiro Type of support:

EMail. Linux installation and configuration, Internet and Intranet conectivity. Sample prices:

Upon request. Last modified:

November 8, 1996.

14 Mexico

14.1 Miguel de Icaza

          Av Copilco 163-22-1003
          Mexico City
          Mexico

Phone:

Mexico-(915)-622-4680 Fax:

None EMail:

miguel.de.icaza@linux.org Url:

None Contact:

Miguel de Icaza Type of support:

Phone, EMail, training. Kernel hacking, driver developement, system administration, applications programming, bug fixing of publicaly available software, extension of freely redistributable software.

Sample prices:

US$30-$50/hour. Last modified:

September 15, 1996.

14.2 Laboratorio de Sistemas Linux

     Casa 1 Manzana 2
     INFONAVIT Lomas de Cortes
     Cuernavaca, Morelos
     Mexico

Phone:

+52+73+170952 Fax:

+52+73+297098 EMail:

turbo@linux.lsl.com.mx Url:

http://www.lsl.com.mx Contact:

Aztec Eagle Turbo, IRL Hector Daniel Cortes Gonzalez Type of support:

EMail, onsite training. General Linux consulting, Linux workstations, networking, Internet connectivity, and web server.

Sample prices:

Upon request. Last modified:

December 1, 1996.

15 The Netherlands

15.1 Bitbybit Information Systems

          Kluyverweg 2A
          2629 HT Delft
          The Netherlands

Phone:

+31.15.2682569 Fax:

+31.15.268.25.30 EMail:

info@bitbybit-is.nl Url:

http://www.bitbybit-is.nl Contact:

Erno Marks Type of support:

Installation, troubleshooting, training, Internet and Intranet connectivity. Sample prices:

US$60-$75/hour. Last modified:

September 13, 1996.

15.2 BitWizard B.V.

          Oosterstraat 23
          2611 TT  Delft
          The Netherlands

Phone:

+31-15-2137459 Fax:

None EMail:

info@BitWizard.nl Url:

http://www.BitWizard.nl Contact:

Rogier Wolff <R.E.Wolff@BitWizard.nl> Type of support:

Troubleshooting, programming, firewall installation, installation, support contracts. Sample prices:

HFL 150/hour. HFL 3000 for a small firewall. Support contracts start at HFL 200/month. Last modified:

September 29, 1996.

15.3 Cistron Internet Services

          PoBox 297
          2400 AG
          The Netherlands

Phone:

+31-172-419445 Fax:

+31-172-430979 EMail:

linux@cistron.nl Url:

http://www.cistron.nl Contact:

Danny ter Haar Type of support:

Author of large parts of Linux (sysvinit, minicom, network code and kernel code). Maintainer of many Debian GNU/Linux packages. Internet and Intranet provider including firewalls and leased lines.

Sample prices:

HFL 150/hour. Last modified:

September 24, 1996.

15.4 MANTRA Research

          Wijenburg 10
          5655 EW Eindhoven
          The Netherlands

Phone:

+31-40-2571950 Fax:

+31-40-2571950 EMail:

naresh@letterbox.com Url:

http://www.focus-asia.com/home/naresh/MANTRAHome Contact:

Naresh Sharma Type of support:

Phone, EMail, inhouse, training. We are a Caldera Channel Partner. Sample prices:

Soon to come on our home page. Last modified:

September 26, 1996.

16 Norway

16.1 it-analyse as

          Breist{\o}lsveien 14
          N-5035 Bergen-Sandviken
          Norway

Phone:

+4755962221 Fax:

None EMail:

magnus@login.eunet.no Url:

None Contact:

Magnus Alvestad Type of support:

Phone, EMail, inhouse, training. Sample prices:

US$80-$100/hour. Last modified:

September 13, 1996.

16.2 NorConnect Internet Services AS

          Boks 9273 Gr\"{o}nland
          N-0134 Oslo
          Norway

Phone:

+47 22 17 41 70 Fax:

+47 22 17 64 99 EMail:

post@norconnect.no Url:

http://www.norconnect.no Contact:

Vidar Hokstad <vidarh@norconnect.no> Type of support:

Phone, EMail, inhouse, training, programming, planning, maintenance of systems, phone support for customers of our customer, porting of programs etc.

Sample prices:

Support agreements starting from 500,- NOK (approximately US$75), a month fixed fee (phone and EMail support, and reduced rates for other services). Projects or one time jobs starting at 650,- NOK (approximately US$110)/hour, with large discounts possible for larger projects, depending on the type of work, and whether our customer wants the project tied to one of our consultants, or not.

Last modified:

September 15, 1996.

17 Pakistan

17.1 Vision Systems

     179-C-1 Model Town
     Lahore-54700
     Pakistan

Phone:

(042) 583-6152 or (042) 586-9353 Fax:

92-42-5834920 EMail:

abnaseem@paknet1.ptc.pk Url:

http://www.freenet.edmonton.ab.ca/˜ abidn Contact:

Abid Nasim Type of support:

Training, configuration and BBS.

Sample prices:

Upon request. Last modified:

September 18, 1996.

18 Poland

18.1 Polskie Sieci Teleinformatyczne S.A.

          ul. Zaolzianska 11
          41-800 Zabrze
          Poland

Phone:

+48 32 176 14 30 Fax:

+48 32 176 01 75 EMail:

pst@netkat.com.pl Url:

http://www.netkat.com.pl Contact:

Andrzej Madloch Type of support:

Phone, EMail, training, web/CGI, networking, Internet connectivity planning, design, instalation, configuration and security.

Sample prices:

US$50-$100/hour dependent on project level. Last modified:

November 25, 1996.

19 South Africa

19.1 Obsidian Systems

          POBox 687
          Honeydew, 2040
          South Africa

Phone:

+27-83-377-4946 Fax:

+27-11-794-1334 EMail:

adw@obsidian.co.za Url:

http://www.obsidian.co.za Contact:

Anton de Wet Type of support:

Onsite, installations, training. Sample prices:

R150/hour, subject to change. Last modified:

November 30, 1996.

20 Sweden

20.1 Guide Datakonsult

          Gustavslundsv. 12
          S-167 51 Bromma
          Sweden

Phone:

+46 70 745 52 39 Fax:

+46 634 46 00 EMail:

mge@guide.se Url:

http://www.guide.se Contact:

Mikael Geijer Type of support:

EMail, inhouse. Sample prices:

On discussion. Last modified:

September 15, 1996.

20.2 Signum Support AB

          Teknikringen 8
          S-583 30 Link\"{o}ping
          Sweden

Phone:

+46 (0)13 21 46 00 Fax:

+46 (0)13 21 47 00 EMail:

info@signum.se Url:

http://www.signum.se Contact:

Magnus Redin <redin@signum.se> Type of support:

Phone, fax, EMail and online. Specialised in the Readynet Internet server and the Readycom conferensing system, both are in house developed Free Software.

Sample prices:

Consulting US$110/hour, Readynet Internet server installation and support approximately US$4300. Last modified:

September 9, 1996.

20.3 Tarkus AB

          Skommartorpsv. 24
          S-735 38 Surahammar
          Sweden

Phone:

(+46) 70 7522246 Fax:

(+46) 220 33907 EMail:

tarkus@tarkus.se Url:

http://www.tarkus.se Contact:

Carlos Vidal, Bo Berggren Type of support:

Mostly via phone or EMail. Inhouse if required. We work in Scandinavia, Germany and Switzerland. Sample prices:

From 0 to 700 SEK/hour (US$0-$100) depending on the job. Many ”easy” questions are answered for free.

Last modified:

September 13, 1996.

21 Turkey

21.1 Hakan Kalyoncu

          Perihan S 49/12
          Sisli 80260 Istanbul
          Turkey

Phone:

+90 532 323 4186 Fax:

None EMail:

hk@yore.com.tr Url:

http://www.yore.com.tr/˜ hk Contact:

Hakan Kalyoncu Type of support:

Phone, EMail, inhouse, onsite. Training, Internet connectivity, Internet server setup, administration. Sample prices:

Free (Personal individual), US$100/hour (Commercial). Last modified:

October 27, 1996.

22 United Kingdom (UK)

22.1 Cygnet Internet Services Ltd

          Lee Valley Technopark
          Ashley Road
          London, N17 9LN
          UK

Phone:

0181 880 4650 Fax:

0181 880 4651 EMail:

steve@cygnet.co.uk Url:

http://www.cygnet.co.uk Contact:

Steve Merrony Type of support:

Phone, EMail, inhouse, training.

Sample prices:

Consultancy from £400/day. Last modified:

November 7, 1996.

22.2 Paul Dunne

          30 Onslow Gardens
          London N10 3JU
          UK

Phone:

+44-(0)181-365-2821 Fax:

None EMail:

paul@tiny1.demon.co.uk Url:

http://www.tiny1.demon.co.uk Contact:

Paul Dunne Type of support:

EMail, also onsite work. Linux installation and configuration, Internet connectivity, system administration, programming with shell and Perl scripts.

Sample prices:

Basic rate £30/hour. Last modified:

October 1, 1996.

22.3 Laboratory for Micro Enterprise

          Oakley Court
          Benson
          Wallingford, OX10 6QH
          UK

Phone:

(44) (0)1491 832932 Fax:

(44) (0)1491 832932 EMail:

lme@storrs.demon.co.uk Url:

http://www.i-way.co.uk/˜ storrs Contact:

John Storrs Type of support:

As required. Our focus is to support very small enterprises (most of us) with affordable systems and tools. See our web pages for details.

Sample prices:

By negotiation. Last modified:

September 15, 1996.

22.4 Nene Information Sytems

          32 Glasgow Street
          Northampton, NN5 5BL
          UK

Phone:

+44 (0)1604 464561 Fax:

+44 (0)1604 464561 EMail:

tim@nisys.demon.co.uk Url:

http://www.nisys.demon.co.uk Contact:

Timothy Peters Type of support:

EMail, inhouse training, software development, network, Internet and connectivity Solutions.

Sample prices:

Upon request. Last modified:

September 26, 1996.

23 United States of America (USA)

23.1 Aegis Information Systems, Inc.

     P.O. Box 730
     Hicksville, NY 11802
     U.S.A.

Phone:

(800) 234-4700 Fax:

(800) 247-1216 EMail:

info@AegisInfoSys.com Url:

http://www.AegisInfoSys.com Contact:

David Harari Type of support:

Phone, EMail, inhouse, training, onsite support, 24 by 7 support, direct online, etc. We also do physical planning and installations, contract programming services, custom EMail forwarding and reverse-forwarding, mid-range and mainframe programming and support, etc.

Sample prices:

US$60-$90/hour for general programming, US$85-$110/hour for special services, US$125-$140/hour for emergency, on-call and round-the-clock support.

Last modified:

October 1, 1996.

23.2 Rahim Azizarab

          1925 S 3rd Ave
          Maywood,  IL  60153
          U.S.A.

Phone:

708-344-6994 Fax:

None EMail:

rahim@megsinet.net or ray@tezcat.com Url:

None Contact:

Rahim Azizarab Type of support:

Phone, EMail, inhouse, training. Sample prices:

Free for individuals. Last modified:

September 18, 1996.

23.3 BRT Technical Services Corporation

          50 Newtown Road
          Danbury, CT 06810
          U.S.A.

Phone:

(203) 748-5100 Fax:

(203) 792-0280 EMail:

info@brttech.com Url:

http://www.brttech.com Contact:

Chad Robinson Type of support:

Phone, EMail, inhouse, training. Sample prices:

Negotiated with client. Last modified:

September 20, 1996.

23.4 Colorado Computer Consultants

          615 Quartz Way
          Broomfield, CO 80020-1727
          U.S.A.

Phone:

303-404-9589 Fax:

303-404-9588 EMail:

co3-info@co3.com Url:

http://www.co3.com Contact:

Robert Kaczanowsk Type of support:

phone, onsite, EMail. Sample prices:

US$35/hourr (individuals), US$50/hour (small business), US$75/hour (corporate). Contract prices available.

Last modified:

October 8, 1996.

23.5 M. Cooper

          PO Box 237
          St. David, AZ 85630
          U.S.A.

Phone:

(520) 720-9431 Fax:

None EMail:

thegrendel@theriver.com Url:

http://personal.riverusers.com/˜ thegrendel Contact:

Leo M. Cooper Type of support:

Phone, EMail. We offer set-up, programming and training. Sample prices:

US$20-$50/hour, negotiable. Last modified:

November 30, 1996.

23.6 Couvares Consulting

          211 West Gilman St., Suite 3W
          Madison, WI 53703
          U.S.A.

Phone:

+1 (608) 256-6901 Fax:

None EMail:

couvares@family.hampshire.edu Url:

None Contact:

Peter Couvares Type of support:

We offer phone/EMail support, installation, ongoing administration, training, programming, and specialized consulting services for Free Software, Linux, and other Unix environments.

Sample prices:

US$50/hour (commercial), US$40/hour (non-profit); sliding scale for personal. Last modified:

September 9, 1996.

23.7 Crynwr Software

          521 Pleasant Valley Rd.
          Potsdam, NY 13676
          U.S.A.

Phone:

+1-315-268-1925 Fax:

+1-315-268-9201 EMail:

info@crynwr.com Url:

http://www.crynwr.com Contact:

Russell Nelson Type of support:

Qualified to do any kind of support; specializing in networking drivers. Sample prices:

US$100/hour; fixed-price quotes for your task. Last modified:

December 1, 1996.

23.8 Daylight Software

          1062 Lexington Lane
          Estes Park, CO 80517
          U.S.A.

Phone:

(970) 586-6062 Fax:

None EMail:

daylight@frii.net Url:

http://www.frii.net/˜ daylight Contact:

Chris Howard Type of support:

Administration and programming via phone and the Internet. Sample prices:

US$65/hour. Last modified:

November 21, 1996.

23.9 DESIGN SCIENCE LABS

          POB 542
          Berea OH 44017-0542
          U.S.A.

Phone:

216/243-1346 Fax:

216/243-1346 EMail:

stutz@dsl.org Url:

http://www.dsl.org Contact:

Michael Stutz Type of support:

Hackers for hire, onsite and remote (phone and Internet). We are exclusive to Linux and GNU and provide comprehensive hypermedia/Internet solutions including WWW design, analog/digital logo/illustration, technical writing (TeX, texinfo, SGML etc) and training.

Sample prices:

Billing by project or hourly. US$50/hour for most work. Last modified:

October 30, 1996.

23.10 DigiTech, Inc.

          301 N. Ankeny Blvd., Suite 100
          Ankeny, IA 50021
          U.S.A.

Phone:

515/964-8032 Fax:

515/965-7626 EMail:

digitech@ecity.net Url:

http://george.ecity.net/˜ digitech Type of support:

Phone, EMail, inhouse, onsite, install, hardware, software, etc. Sample prices:

          Hardware/software support                     US\$52/hour
          Networking                                    US\$65/hour

Last modified:

September 15, 1996.

23.11 Electronic Information Systems

     8758 Brook Road
     McLean, VA 22102
     U.S.A.

Phone:

(703)281-0409

Fax:

None

EMail:

ghurwitz@tjhsst.edu

Url:

http://bhurwitz.erols.com/SS

Contact:

Gus Hurwitz

Type of support:

EMail preferred, phone very doable, inhouse can be arranged, individual training with inhouse, group training available also.

Sample prices:

Simple questions via EMail are generally handled at no charge, ongoing EMail consulting US$25-$50, ongoing EMail support contract (seperate mail addresses available, are handeled first) available for about US$10 a month for 6 months or more. Phone support available if local to (703)281 exchange, or you pick up the tab, generally used as an extension to EMail support. Inhouse setup as little as US$50, with network setup about US$100, if hardware setup about US$150. The more systems we setup, the less it costs. Classes available for more than 5 people: topics ranging from simple installation, to C/C++ development, to system administration, to server setup and administration, etc. Students are generally welcome (and encouraged) to maintain EMail corespondance after the class, asking questions, and sharing experiences. If hourly rates are perfered, we like to stay below US$30, and are often willing to work for advertising space or endorsements.

Last modified:

September 20, 1996.

23.12 Etc Services

     31 N. Vivyen St.,
     Bergenfield, NJ 07621-1529
     U.S.A.

Phone:

(201) 385-7113 Fax:

None EMail:

swebster@carroll.com Url:

None Contact:

Scott D. Webster Type of support:

Inhouse. System setup and administration, TCP/IP network configuration, web server setup (not content creation), Linux LAN servers, DNS configuration, experience with Linux on x86, Linux on DEC Alpha (AXP), Digital Unix on Alpha (AXP), and Solaris on SPARC and UltraSPARC.

Sample prices:

Call to negotiate. Last modified:

September 15, 1996.

23.13 Evolve Computer Solutions

     1523 E. Valley Parkway, Suite 237
     Escondido, CA 92027
     U.S.A.

Phone:

(619)-746-2757 Fax:

(619)-751-0612 EMail:

evolve@sierra.net Url:

None Contact:

Sean P. Kane Type of support:

Phone, EMail, inhouse, training. Sample prices:

US$40/hour. Last modified:

September 17, 1996.

23.14 Christopher J. Fearnley

          4807 Chester AVE APT 207
          Philadelphia PA 19143-3441
          U.S.A.

Phone:

215-724-2265 Fax:

None EMail:

cjf@netaxs.com Url:

http://www.netaxs.com/˜ cjf Contact:

Chris Fearnley Type of support:

Phone, inhouse, Internet. Sample prices:

US$500/day. Last modified:

September 17, 1996.

23.15 Granite Computing Solutions

          P.O. Box 270103
          Fort Collins, CO 80527-0103
          U.S.A.

Phone:

+1 (970) 225-2370 Fax:

None EMail:

granite@SoftHome.Net Url:

http://www.SoftHome.Net/granite Contact:

Brian Grossman <brian@SoftHome.Net> Type of support:

In person, phone, EMail, Internet remote (telnet), training developed on request for Internet and Unix topics.

Sample prices:

US$45/hour for technical support. Last modified:

September 14, 1996.

23.16 Steve Hedlund

          Las Vegas, NV
          U.S.A.

Phone:

702-254-6749 Fax:

None EMail:

ssh@nevadacom.com Url:

http://nevadacom.com Contact:

Steve Hedlund Type of support:

Phone, EMail, inhouse, training. Sample prices:

US$25-$40/hour. Last modified:

September 12, 1996.

23.17 I-Link, Inc.

          136 Forests Edge Place,
          Laurel, MD 20724
          U.S.A.

Phone:

301-604-8411 Fax:

301-604-8411 EMail:

fbennett@i-linkcom.com Url:

http://i-linkcom.com Contact:

Fred Bennett Type of support:

Phone, EMail, inhouse, training. We support software from Red Hat Software, InfoMagic, Walnut Creek, and Caldera.

Sample prices:

US$50-$75 Hour. Last modified:

October 12, 1996.

23.18 InfoMagic, Inc

     11950 N. Highway 89
     Flagstaff, AZ 86004
     U.S.A.

Phone:

1-900-786-5555 Fax:

1-520-526-9573 EMail:

support@infomagic.com Url:

http://www.infomagic.com/support/linux Contact:

Henry M. Pierce Type of support:

Individuals who purchase InfoMagic Linux Developer’s Resource get free EMail and fax back support. Support via phone for our Linux Developer’s Resource or any Linux vendor’s Linux CD is US$2.00/minute in the U.S.A. Everyone is welcome to look at our support web pages. International customers needing phone support should contact us via EMail first.

Sample prices:

Fixed service contracts, including remote maintance are available starting at US$50/engineering hour. Please contact us by EMail for additional information.

Last modified:

September 28, 1996.

23.19 Internet Robotics

     131 Willow Lane, Floor 2
     Oxford, OH 45056
     U.S.A.

Phone:

513.523.5028 Fax:

513.523.7501 EMail:

support@bill.iac.net Url:

http://www.iac.net/˜ mikeh Contact:

Mike Heins Type of support:

Specializing in online shopping with MiniVend, the GPL shopping cart. Expertise in all aspects of Linux, WWW administration, shell and Perl scripting and network configuration. Performance tuning. Phone and EMail, will travel for training courses.

Sample prices:

          Expert technical support (uncontracted):      US\$100/hour
          Monthly retainer (includes first hour):       US\$50
          Contract technical support (with retainer):   US\$85/hour
          Contract shopping software design:            US\$85/hour
          Contract shopping site HTML design:           US\$50/hour
          Configuration support:                        US\$50/hour

Last modified:

September 12, 1996.

23.20 Internet Systems and Services

          13628 Bent Tree Circle, Suite 201
          Centreville, VA 22020
          U.S.A.

Phone:

(703) 222-4243 Fax:

(703) 222-7320 EMail:

bass@silkroad.com Url:

http://www.silkroad.com Contact:

Tim Bass, Principal Systems Engineer Type of support:

Phone, EMail, inhouse. Sample prices:

US$85.00/hour plus travel expenses (20 hours normal minimum). Last modified:

September 15, 1996.

23.21 Isle-Net Telecommunications

          PO Box 434
          Port Monmouth, NJ 07758
          U.S.A.

Phone:

(908) 787-0663 Fax:

None EMail:

dan@islenet.com Url:

None Contact:

Dan Srebnick Type of support:

Specializing in providing efficient, low-cost Internet connectivity solutions for business and government, using Linux.

Sample prices:

Hourly consulting or turnkey package solutions are available. Last modified:

October 12, 1996.

23.22 Matthew Kaylor

          13639 Barnhart Rd.
          Clear Spring, MD 21722
          U.S.A.

Phone:

(301) 733-2922 Fax:

None EMail:

maxit@nfis.com Url:

http://maxit.nfis.com Contact:

Matthew Kaylor Type of support:

EMail, IRC, phone, world wide onsite support services. Linux system setup (remote or in house) to commercial networked Internet servers (complete). Personal or group Linux training courses.

Sample prices:

          Commercial - US\$5000/month, US\$150/day, US\$30/hour
          Personal Individual - Based on needs

Last modified:

October 7, 1996.

23.23 Kensho, LLC

          9444 Balboa Avenue Suite 235
          San Diego, California 92123
          U.S.A.

Phone:

+1 (619) 514-3939 Fax:

+1 (619) 514-3938 EMail:

info@kensho.com Url:

http://www.kensho.com Contact:

Sean Dreilinger <sean@kensho.com> Type of support:

Research, training, and marketing of networked information services, with an eye towards PD and GPL solutions, i.e. Linux!

Sample prices:

Free consultation. Per project and hourly rates negotiable. Last modified:

September 12, 1996.

23.24 Kracked Rock Komputing, LLC

     13 Fleetwood Drive
     Danbury, CT 06810-7009
     U.S.A.

Phone:

+1 203 778 6269 Fax:

+1 203 778 6292 EMail:

ken@eci.com Url:

None Contact:

Kenneth E. Nawyn Type of support:

Phone, EMail, inhouse systems administration, system configuration and tech support for Linux, FreeBSD, SunOS, Solaris, IRIX and AIX. Over ten years of Unix systems adminstration experience along with Perl and C coding experience. For more detailed description of our experience and services, please contact us.

Sample prices:

Base rate US$100/hour. For a more detailed rate schedule, please contact us. Last modified:

September 24, 1996.

23.25 Kymsoft Consulting

          745 Foothill Dr
          Eugene, OR 97405
          U.S.A.

Phone:

541-465-9008 Fax:

541-465-9052 EMail:

rick@efn.org Url:

http://www.efn.org/˜ rick Contact:

Rick Bronson Type of support:

Phone, EMail, inhouse, training (12 years Unix experience). Sample prices:

US$65/hour. Last modified:

September 12, 1996.

23.26 LANscape

          1100 W Lloyd Expressway
          Evansville, IN 47708
          U.S.A.

Phone:

(800) 896-6171 or (812) 423-1231 Fax:

(812)426-6138 EMail:

info@lanscape.net Url:

http://www.lanscape.net Contact:

Steven Pritchard Type of support:

Limited free phone and EMail support available for Linux users as time permits. Ongoing phone and/or EMail support available with a service contract. TCP/IP network setup and maintenance, system administration, programming, and training available on-site or at/from our office. We also offer web site hosting, POP mail accounts, EMail forwarding, primary and secondary DNS, etc. See our web site for details.

Sample prices:

Base rate is US$100/hour for most services. Contract pricing is available at a reduced rate. Last modified:

October 22, 1996.

23.27 LOD Communications

     1621 Creighton Road
     Pensacola, FL 32504
     U.S.A.

Phone:

904.494.0955 Fax:

N/A EMail:

support@lod.com Url:

http://www.lod.com/cns.html Contact:

Todd Lawrence <lawrence@lod.com> Type of support:

We provide all forms of Linux/Unix consultation, phone, telnet/rlogin, and onsite. We provide training in all forms of Unix system administration and networking. We can generally provide our own transportation anywhere in the continental USA for onsite consultation.

Sample prices:

     $30 - Phone Support
     $50/hour with fixed/flat rate estimates for remote (telnet/rlogin) consulting
     $75/hour with fixed/flat rate estimates for onsite Unix consultation

Last modified:

October 18, 1996.

23.28 NDA

          400 West Cummings Park
          Woburn, MA 01801
          U.S.A.

Phone:

617-937-3338 Fax:

617-937-3775 EMail:

info@nda.com Url:

http://www.nda.com Contact:

Tom Bills, Mary Ede Type of support:

Onsite, remote, EMail, phone, training. Sample prices:

US$50-$150 based on task. Last modified:

November 24, 1996.

23.29 NET-Community

          38403 Pelton Road
          Willoughby, OH 44094
          U.S.A.

Phone:

1-800-919-0060 Fax:

1-216-953-5829 EMail:

sales@net-community.com Url:

http://www.net-community.com Contact:

Scott Christley Type of support:

Phone, EMail. NET-Community specializes in support for GNUstep and MediaBook. Our web site has details regarding the support plans.

Sample prices:

          US\$50/year for GNUstep Developer - EMail only
          US\$200/year for GNUstep Corporate - EMail only
          US\$2500/year for MediaBook Comprehensive - EMail and phone

Last modified:

September 20, 1996.

23.30 Neu-Direct Inc.

          8121 Old York Road. Suite 200
          Elkins Park, PA 19027
          U.S.A.

Phone:

215-886-4200 Fax:

215-886-5300 EMail:

pdooley@neu-direct.com Url:

http://www.neu-direct.com Contact:

Patrick Dooley Type of support:

EMail, inhouse, onsite, custom machines.

Sample prices:

US$95/hour onsite, US$60/hour inhouse, and system prices varies on configuration. Last modified:

October 24, 1996.

23.31 The Network Wizards

          2275 West Birdsong Rd
          West Valley City, UT 84119
          U.S.A.

Phone:

(801) 973-4170 Fax:

(801) 975-7355 EMail:

tnw@ix.netcom.net Url:

None Contact:

Pat Eyler <pate@gnu.ai.mit.edu> Type of support:

Phone or EMail (most things); inhouse, training, contract telework available. Sample prices:

Interesting work and worthy charities are free (my definition of worthy). Others average US$50/hour (corporate) or US$25/hour (personal).

Last modified:

September 12, 1996.

23.32 Alvin Oga

          RING, Inc
          P.O. Box 70958
          Sunnuyvale, CA 94086
          U.S.A.

Phone:

408-245-8400 Fax:

408-523-9097 EMail:

alvin@planet.fef.com Url:

http://planet.fef.com or http://www.linux-consulting.com Contact:

Alvin Oga Type of support:

EMail, and over the Internet. We install, setup, test and maintain your Linux machines and Internet connectivity networking.

Sample prices:

Upon request. Last modified:

November 5, 1996.

23.33 onShore, Inc.

     1655 E. 55th St. \#1
     Chicago, IL 60615
     U.S.A.

Phone:

(312) 667-4673 Fax:

(312) 667-4263 EMail:

sales@onshore.com Url:

http://www.onshore.com Contact:

Stelios Valavanis Type of support:

Inhouse installaion and support of Linux for custom solutions such as routing, firewall, EMail, web serving, etc. Custom solutions for job tracking, print spooling/processing, Intranet, etc.

Sample prices:

Support hourly rates are US$110/hour standard with lower rates for pre-paid blocks. Development rate is US$125/hour.

Last modified:

October 9, 1996.

23.34 Practical Network Design

     9 Chambers Rd
     Mansfield, OH 44906-1302
     U.S.A.

Phone:

(419) 529-3841 Fax:

(419) 529-3625 EMail:

stick@richnet.net Url:

None Contact:

Chuck Stickelman Type of support:

Inhouse, phone, EMail. Can provide system and network planning, installation, support and administration. Training is also available for hardware service/support, system/network administration, and various user-level applications.

Sample prices:

Phone support is US$50/hour (1 hour minimum). Inhouse support can be contracted on an hourly basis at US$100/hour or by day/half-day or by project.

Last modified:

September 23, 1996.

23.35 Probe Technology Inc

     10159 J St
     Omaha, NE 68164
     U.S.A.

Phone:

402-593-9800 Fax:

402-593-8748 EMail:

info@probe.net Url:

http://www.probe.net Contact:

Dale Botkin, Tim Russell Type of support:

Phone, EMail, training, configuration. Can configure Internet routers, firewalls, mail servers, web servers (Apache), etc. Can also provide hardware platform of choice at competitive prices, as well as provide a near-turnkey solution for an ISP setup. Training available.

Sample prices:

US$100/hour consulting, average setup of new system with services is 5-7 hours. Last modified:

October 9, 1996.

23.36 PSA AnswerDesk – Pencom Systems Administration

     9050 Capital of Texas Hwy North, Suite 310
     Austin, TX 78759
     U.S.A.

Phone:

(512) 343-1111 or contact your nearest Pencom office. Fax:

(512) 346-6444 EMail:

answer-info@pencom.com Url:

http://www.pencom.com/psa/answer.html Contact:

Randy Kirchhof Type of support:

AnswerDesk is an Open Systems phone support desk specifically designed to provide a backstop with difficult, senior administration issues. We do not provide end user support. We support administrators working with all Unix platforms, including Linux. Our client list includes a wide variety of Fortune 500 companies utilizing open systems, many with mission critical support requirements. Our technical database is supported by 200 of the best systems administrators in the business.

Sample prices:

AnswerDesk charges US$100/hour, with a one hour minimum. If we don’t solve your problem, you are not charged. Please request a brochure for more in-depth info.

Last modified:

September 24, 1996.

23.37 RCT Design

     Santa Barbara, CA 93117
     U.S.A.

Phone:

805-683-9717 Fax:

805-683-0717 EMail:

bob@zooid.com Url:

http://www.zooid.com Contact:

Bob Tellefson Type of support:

Turnkey Linux Systems with support, Internet servers, web site design and CGI programming, remote system maintenance and/or training by phone or EMail.

Sample prices:

US$45/hour. Last modified:

November 30, 1996.

23.38 Reboot Inc.

     748 N Lindberg \#180
     St. Louis, MO 63031
     U.S.A.

Phone:

314-731-1034 Fax:

314-731-1034 EMail:

reboot@i1.net Url:

http://www.i1.net/˜ reboot Contact:

Zubin Chandran Type of support:

Onsite support to the 314 and 618 area codes, phone support to other areas. Training onsite as well as hardware upgrades and sales. Will also handle EMail requests.

Sample prices:

US$27 for up to the first hour and 15 minutes, then US$5.50 every 15 minutes thereafter for onsite support. Phone support is 60 cents per minute. Call for special pricing and promotional deals.

Last modified:

November 25, 1996.

23.39 RG Consulting

     1751 East Roseville Pkwy. \#1828
     Roseville, CA 95661
     U.S.A.

Phone:

+1 916 786 7945 Fax:

+1 916 786 5311 EMail:

rfg @ monkeys . com (Please note that this is no correct address; you have to remove the blanks before you can send messages to it. Sorry for the inconvenience – this is against spammers)

Url:

None Contact:

Ronald F. Guilmette Type of support:

Providing offsite and onsite support, maintenance, and enhancement of GNU software development tools.

Sample prices:

Upon request. Last modified:

October 10, 1996.

23.40 Bill Rousseau

     842 Lucille Street
     Livermore, CA 94550-3519
     U.S.A.

Phone:

510/455-8008 Fax:

510/455-8008 (usually it works) EMail:

rousseau@aimnet.com Url:

Look me up in http://www.ieee-sv-consult.org which I maintain. Contact:

Bill Rousseau Type of support:

Support available onsite, via phone, or EMail. Software development is offsite as much as possible. Training is available if there is sufficient interest.

Sample prices:

General computing consulting (Unix, Linux, network configuring, software development) US$100/hour. Support for individuals attempting to install or configure Linux for their own use or for work that promises to benefit the Linux community may be offered at much lower prices or for free in special cases.

Last modified:

October 18, 1996.

23.41 Softcraft Impresa

     2923 Cohansey Drive
     San Jose, California 95132
     U.S.A.

Phone:

+1.408.251.9820 Fax:

None EMail:

mauro@softcraft.com Url:

http://www.softcraft.com Contact:

Mauro DePalma Type of support:

We can fine tune homogeneous and heterogeneous networks with Linux/Unix nodes and others TCP/IP capable. We also have an evolving secure operating environment based on Linux which is complete but does not (yet) come with installation scripts/programs and is currently only available for Intel. We’ll call it Linux generations. Gladly demo it and use it to build a LAN/WAN with shadow/NYS/DNS for secure nodes, gateways and networks.

Sample prices:

US$90/hour. Depends on the kind of work and duration.

Last modified:

September 20, 1996.

23.42 Sophic Consulting

     4822 Wrenwood Place
     Columbus, IN
     U.S.A.

Phone:

(812) 342-0166 Fax:

(812) 342-0166 EMail:

sophic@mail.cpbx.net Url:

http://www.sophic.com Contact:

Bruce Varney, Chief Technical Officer Type of support:

Phone, EMail, inhouse, onsite. Custom software development, system configuration and administration support for Linux, cross-platform user interface development, Linux Internet/web servers, web page development.

Sample prices:

Pricing on a contractual per-project basis, but based on targets of US$25-$100/hour depending on the project.

Last modified:

October 11, 1996.

23.43 Earl A. Stutes, Inc.

     P.O. Box 35567
     Monte Sereno, CA 95030-0567
     U.S.A.

Phone:

408.448.1089 Fax:

408.264.5129 EMail:

estutes@eas.com Url:

http://www.eas.com Contact:

Earl Stutes Type of support:

Phone, EMail, inhouse, training. Sample prices:

Upon Request. Last modified:

September 12, 1996.

23.44 Triangle Software Corporation

          3216 Caldwell Road
          Atlanta, GA 30319
          U.S.A.

Phone:

1-800-625-7787, 770-455-9089 Fax:

Upon request. EMail:

triangle@america.net Url:

None Contact:

Ben Cooper Type of support:

Inhouse ANSI C++ training with labs using G++ (very well reviewed by students).

Sample prices:

US$5000 plus travel expenses for a five-day C++ class with up to 15 students, taught at your site. Last modified:

December 1, 1996.

23.45 tummy.com, ltd.

          300 South 16th Street, Suite 807
          Omaha, NE 68102
          U.S.A.

Phone:

402-344-4426 Fax:

402-341-7119 EMail:

info@tummy.com Url:

http://www.tummy.com Contact:

Sean Reifschneider <jafo@tummy.com> Type of support:

Phone, EMail, inhouse. Custom programming, firewall setup, general Unix/Linux consulting. Sample prices:

Depending on level of expertise, usually in the US$60-$125 range. Last modified:

October 1, 1996.

23.46 Vennerable Consultants

          7216 Lincoln Avenue, Suite 101
          El Cerrito, CA 94530 3240
          U.S.A.

Phone:

510 525 2989 Fax:

510 525 8072 EMail:

jason@vennerable.com Url:

http://www.vennerable.com Contact:

Jason Venner Type of support:

Phone, EMail, network and inhouse. Sample prices:

Variable by job and customer. Last modified:

September 10, 1996.

23.47 Vertex Development

          11 Crestview Court
          Potomac, MD 20854
          U.S.A.

Phone:

301-340-2162 Fax:

301-309-6968 EMail:

jeff@vertexdev.com Url:

http://www.vertexdev.com Contact:

Jeff Barr Type of support:

Porting Unix and non-Unix software to Linux. User Interface design, development, and review. Applications and Systems software development.

Sample prices:

Porting and development projects are typically handled on a firm-fixed price basis. Other services are supplied at US$120-$150/hour, depending on duration and travel requirements.

Last modified:

September 15, 1996.

23.48 Gregg Weber

          1076 Carol Lane \#3
          Lafayette, CA 94549
          U.S.A.

Phone:

510 283-6264 Fax:

None EMail:

gregg@netcom.com Url:

None Contact:

Gregg Weber Type of support:

On site in SF bay area, or can travel for big jobs. Sample prices:

US$55/hour. Last modified:

September 9, 1996.

23.49 Jim Willette

          18190 - 15 Mile Rd
          Big Rapids, MI 49307
          U.S.A.

Phone:

(616) 592-3119 Fax:

None EMail:

jimw@ctec.net Url:

http://www.netonecom.net/˜ jimw, http://www.ctec.net Contact:

Jim Willette Type of support:

Phone, EMail, inhouse, telnet. Specialties include Internet firewalls, proxy servers, web servers, domain name servers (DNS), Linux terminal servers, Internet connectivity networking, NFS and more.

Sample prices:

US$40/hour phone support. US$50/hour onsite plus expenses and travel time (negotiable). Last modified:

October 13, 1996.

23.50 Worldwide Solutions, Inc.

     4450 Arapahoe Ave., Suite 100
     Boulder, CO 80303
     U.S.A.

Phone:

+1.303.581.0800 Fax:

None EMail:

info@wwsi.com Url:

http://www.wwsi.com Contact:

Steve Hultquist and Blaine Berger Type of support:

Phone, EMail, Internet, onsite, training, installations, configuration, specification, architecture, engineering, integration, etc.

Sample prices:

Variable. Monthly support agreements or project based statement of work pricing preferred. Standard hourly rates available for those with specific requirements.

Last modified:

December 1, 1996.

23.51 xprt Computer Consuting, Inc.

     731 Voyager
     Houston, TX 77062
     U.S.A.

Phone:

713 480 8649 Fax:

713 480 7357 EMail:

jody@sccsi.com Url:

http://zeus.xprt.com/˜ info/xprt.html Contact:

Jody Winston Type of support:

Phone, EMail, training, custom software, custom device drivers, custom application development. Sample prices:

Either by the hour starting at US$100 or a firm fixed price. Last modified:

September 12, 1996.

23.52 Yggdrasil Computing, Inc.

     4880 Stevens Creek Boulevard, Suite 205
     San Jose, CA 95129-1034
     U.S.A

Phone:

(800) 261 6630 Fax:

(408) 261 6631 EMail:

tech-support@yggdrasil.com Url:

http://www.yggdrasil.com Contact:

Patrick Bedell Type of support:

We provide phone and EMail support. We are also able to use our T1 Internet connection to solve your problem online (if your machine is also on the Internet).

Sample prices:

US$25 for a 15 minute support call guaranteed to solve your problem. Other arrangements are available.

Last modified:

September 27, 1996.

23.53 ZEI Software

     2713 State Hwy 23 South
     Newfoundland, NJ 07435
     U.S.A.

Phone:

201-208-8800 Fax:

201-208-1888 EMail:

art@zei.com Url:

None Contact:

Arthur Zysk Type of support:

Phone, EMail, inhouse, onsite, training. Sample prices:

Upon request. Last modified:

October 18, 1996.